OpenStudy (dan815):

Anyone know matlab, How do i generate a random positive1000 values that are normally distributed and have a mean of 1 and a standard deviation of 0.27: What I have so far randn(1000,1) this function gives 1000 normally distributed points in 1 column it has a preset standard deviation of 1 and a mean of 0 so if i do randn(1000,1) +1 this gets my mean to 1, however STD is still 1, is there another built in function or wanna come up with another way of generating these points ? There is the normal distrubtion equation given by \[f(x)=1/s\sqrt{2*pi}\exp{-(-x-x_bar)^2/s^2}\]

OpenStudy (dan815):

\[f(x)=\frac{1}{s\sqrt{2*pi}}*e^{{\frac{-(-x-x_{bar})^2}{s^2}}}\]

OpenStudy (dan815):

i really dont want to try to think of generating this data set manually with this equation, is there some built in functions and parameters im missing or possibly a simple way to recode that built in function randn such that the deviation is 0.27 instead of 1

OpenStudy (dan815):

if i multiplying the whole data set by 0.27, then the standard deviation is 0.27 however the mean is also 0.27 now

OpenStudy (dan815):

oh :O

OpenStudy (dan815):

i see lol i got it!